Database last updated on Mon, 02 Feb 2026 06:40:53 +0000 by localhost

Instrument constant changes for the week before 2024051 (1 sec, 0 rows)

HG burned out? (AP results when HG is ON) (0 sec, 13 rows)

Measurements and tests summary.


_jday_loc_serial_skc
051STONY PLAIN#195menu
(0 sec)

Dead time (0 sec)

Nothing to report.

Standard lamp R6 (1 sec)

Nothing to report.

Standard lamp R5 (1 sec)

Nothing to report.

Number of days since last dome/window cleaning (0 sec, 2 rows)

serialdays
195more than a year

Missing files (0 sec, 1 rows)

Seriallast datadays ago
19520240520-1

Run/Stop results for last week. (0 sec, 1 rows)

SerialTotalFailed 0Failed darkFailed 1Failed 2Failed 3Failed 4Failed 5Failed double
195700000000

Ozone results for good DS obs (std<2.5DU, mu<4) within 3 minutes between 2 Brewers. (1 sec)

No ozone measurements to display.

Ozone results for good FM obs (std<5DU, mu<4) within 3 minutes between 2 Brewers. (0 sec)

No ozone measurements to display.

Unsorted comments (1 sec, 12299 rows)

SerialtypeTextCountYYYYJJJ
195env:DosBox12024051
195cubdsp:112024051

The END (0 sec, 2 rows)

If you have questions and/or comments about this page please contact Volodya Savastiouk